OpenStudy (idealist10):

Solve the Bernoulli equation y'+y=y^2.

OpenStudy (idealist10):

Here's the work: n=2 v=1/y y=1/v -1/v^2*dv/dx+1/v=1/v^2 dv/dx-v=-1 u(x)=e^-x -->the integrating factor (e^-x)v'-ve^-x=-e^-x e^-x*v=e^-x+C v=(e^-x+C)/(e^-x) 1/y=(e^-x+C)/(e^-x) y=e^(-x)/(e^(-x)+C) But that's not the answer. The answer is y=1/(1-ce^x) in the book. What's wrong?

OpenStudy (anonymous):

No, you're right. That *is* the answer. Try multiplying by \(\dfrac{e^x}{e^x}\).

OpenStudy (idealist10):

Wait a minute.

OpenStudy (idealist10):

I got 1/(1+ce^x) instead of 1/(1-ce^x).

OpenStudy (anonymous):

Well \(C\) is an arbitrary constant. I don't think it matters. Or maybe it's a typo. Wolfram Alpha has the same answer as you: http://www.wolframalpha.com/input/?i=y%27%2By%3Dy%5E2

OpenStudy (idealist10):

Does anyone know the input of this problem on Ti-89?

OpenStudy (anonymous):

Refer to the attachment from WolframAlpha.

OpenStudy (idealist10):

Thank you guys for the help.